A 75 W light bulb is used for 9 hours a day. The electricity for this bulb comes from a natural gas fired power plant that operates with 46% efficiency. How much CO2 is emitted per day to power this light bulb? Make a simple estimate of the CO2 emissions from the power plant, ignoring transmission losses.
Answer #1

In the question here it is asked to find the value of Carbon dioxide gas emitted in a day to power the light bulb of 75 Watts which runs for 9 hours. Let's solve the problem in steps and find the answer.

We first find the total energy in KWh needed by the bulb in one year = \frac{75}{1000}\times 9\times 365 = 246.375 KWh

Thermal energy content of natural gas = 13,859.8 KWh/ton

But the plant runs with 45% efficiency, hence the thermal energy content of the plant is = = 13,859.8\times 0.45 KWh/ton = 6236.91 KWh/ton

1 ton = 907.185 Kg

Therefore, the thermal energy content of the plant in KWh/Kg is equal to = \frac{6236.91}{907.185} = 6.875 KWh/Kg

To power 75 watts of bulb the amount of natural gas burned was = \frac{246.375}{6.875} = 35.836 Kg

So, 35.836 Kg of natural gas was burned in one year to power the light bulb for 9 hours.

Amount of natural gas burned for one day = \frac{35.836}{365} = 0.098 Kg

Therefore, 0.098 Kg of natural gas burned for one day.

Now, we know that by international conventions the amount of CO2 emitted to in burning 1 Kg of natural gas = 0.2 Kg of CO2

Therefore CO2 emitted in one day = 0.098\times0.2 = 0.0196 Kg = 19.6 grams of CO2

Therefore, we found that 19.6 grams of Carbon dioxide gas is emitted to power 75 watt of bulb for 9 hours continuously.